OpenStudy (anonymous):

y varies indirectly as twice x. When x = 9, y = 2. Find y when x = 3.

OpenStudy (anonymous):

From what i understand its saying that every time x gains a value of 1, y gains a value of 2, therefore the slope of the line would be 2 From the given point and the slope we can determine the y-intercept 2 =2(9) + b 2-18 = b b= -16 Now we have the standard equation of a line \[y = mx +b\] We can plug in the x given, the m and the b we found to get your answer y = 2(3) - 16 y = 6-16 y = -10

OpenStudy (agentnao):

There is a formula for indirect variation (aka inverse variation) and that is y = k over x But since your question says "twice x", it would become y = k over 2x Find k first by using 9 and 2 Once you have found k, you can now find y when x = 2 by plugging them in again.

OpenStudy (agentnao):

2 = k/2(9) 2 = k/18 36 = k y = 36/2(3) y = 36/6 y = 6 when x is 3